OpenStudy (liizzyliizz):

For which of these functions f(x) does the limit f(x) = 2 as x -> -∞? a) (x-2)/(3x-5) b) (2x)/radical(x-2) c) (2x^2-6x+1)/(1+x^2) d) (2x-1)/(x^2+1) e) none of the above For which of these functions f(x) does the limit f(x) = 2 as x -> -∞? a) (x-2)/(3x-5) b) (2x)/radical(x-2) c) (2x^2-6x+1)/(1+x^2) d) (2x-1)/(x^2+1) e) none of the above @Mathematics

myininaya (myininaya):

(a) =1/3 (degrees are "samesy") (b) DNE (sqrt(-number) is imaginary) (c)=2/1=2 (degrees are "samesys") (d)=0 (degree on bottom>degree on top)

OpenStudy (liizzyliizz):

Oh boy I can't believe I forgot all of this. Care to help me refresh my memory so I can not feel like such a failure?